OpenStudy (anonymous):

I am trying to factor by removing the GCF and checking but I cant seem to get it right I have: 3x squared+5xy-10yx squared

OpenStudy (cwrw238):

gcf = x

OpenStudy (anonymous):

\[3x^2+5xy-10yx^2 = x(3x+5y-10yx)\] Therefore, gcf = x